OpenStudy (anonymous):

Find all the solutions of the equation. Leave answers in trigonometric form. X^4 + 625 = 0

OpenStudy (anonymous):

@Koikkara @johnweldon1993

OpenStudy (anonymous):

First lets right this in a better from, no? x^4 = -625

OpenStudy (anonymous):

We want to find all the trigonometric roots, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

For instance x^3 - 27 = 0 so we got r=3 but for this idk how we get r....

OpenStudy (anonymous):

im thinking hmmm... I believe... -625 = 625(cos(pi(1 + 2k)) + i sin(pi(1 + 2k))) (thinking more... I truly hate trig sometimes)

OpenStudy (anonymous):

The number 27 is the solution for x^3 - 27= 0

OpenStudy (anonymous):

i meant 28

OpenStudy (anonymous):

but for -625 idk how to get r

zepdrix (zepdrix):

Oh oh, converting to i like that was silly. Just change -1 to sines and cosines, maybe that will make more sense.\[\Large\rm x=\sqrt[4]{-625}\]\[\Large\rm x=\pm 5 \sqrt[4]{\color{orangered}{-1}}\]So we can rewrite our -1 in terms of sine and cosine,\[\Large\rm x=\pm 5 \sqrt[4]{\color{orangered}{\cos \pi+\mathcal i \sin \pi}}\]